CTL (Concrete Testing Lab) borrowed $120,000 for new equipment at 12% per year, compounded quarterly. It is to be paid back over 4 years in equal quarterly payments.


How much interest is in the 6th payment? $



How much principal is in the 6th payment? $



What principal is owed immediately following the 6th payment?
Answer #1

(a) Interest in the 6th payment is $ 2,651.79

(b) Principal in the 6th payment is $ 6,901.51

(c) Principal owed immediately following the 6th payment is $ 81,491.60

Step-1:Calculation of quarterly payment
Quarterly payment = Loan amount / Present Value of annuity of 1
= $ 1,20,000.00 / 12.561102
= $       9,553.30
Working:
Present Value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.03)^-16)/0.03 i = 12%/4 = 0.03
= 12.56110203 n = 4*4 = 16
Step-2:Loan amortization Schedule
Quarter ended Loan amount Interest Expense Quarterly payment Reduction of principal Ending Loan amount
a b=a*0.03 c d=c-b e=a-d
1 $ 1,20,000.00 $       3,600.00 $ 9,553.30 $ 5,953.30 $ 1,14,046.70
2 $ 1,14,046.70 $       3,421.40 $ 9,553.30 $ 6,131.90 $ 1,07,914.80
3 $ 1,07,914.80 $       3,237.44 $ 9,553.30 $ 6,315.86 $ 1,01,598.94
4 $ 1,01,598.94 $       3,047.97 $ 9,553.30 $ 6,505.33 $     95,093.61
5 $     95,093.61 $       2,852.81 $ 9,553.30 $ 6,700.49 $     88,393.11
6 $     88,393.11 $       2,651.79 $ 9,553.30 $ 6,901.51 $     81,491.60